OpenStudy (anonymous):

A recipe that makes 6 servings calls for 1/2 cups of orange juice and 5/6 cup of lemonade. Nancy modifies the recipe so that it can serve 8. How many cups of lemonade does Nancy need?

OpenStudy (kropot72):

First find the amount of lemonade needed for 1 serving by dividing 5/6 by 6. \[\large Lemonade\ for\ 1\ serving=\frac{\frac{5}{6} }{6}=\frac{5}{6\times6}\] Now multiply the amount of lemonade for 1 serving by 8: \[\large Lemonade\ for\ 8\ servings=\frac{5\times8}{6\times6}=(you\ can\ calculate)\]